OpenStudy (anonymous):

what is the derivative of y=4e^3x^2

OpenStudy (anonymous):

ok, with this problem you need to know the product rule Product rule: F'S+FS' F=4 F'=0 S=e^(3x^2) S'=6x*e^(3x^2) (FYI, whenever you e^(a number) is always e^(a number), in other words, when you derive base e to a power of something, it is always the same. derive of e^2 is e^2. always stay the same. However, e^(3x^2) = 6xe^(3x^2) ^ notice how i derive the exponent and put it in front of e and the expo still stays the same. now to put everything together. (0)(e^(3x^2) + (4)(6xe^(3x^2) = your final answer 0 + 24xe^(3x^2)